OpenStudy (anonymous):

A woman rides a bicycle for 3 hours and travels 36 kilometers. Find the angular velocity of the wheel if the radius is 30 centimeters

OpenStudy (anonymous):

r=.3m C=2pir=.6pi m angvelocity=rev/sec=rad/sec 36000m/(0.6pi m)=#rev t=3(3600)s

OpenStudy (anonymous):

can u plz clarify